OpenStudy (anonymous):

Each side of a square is increasing at a rate of 7cm/s. At what rate is the area of the square increasing when the area of the square is 64cm^2

OpenStudy (bahrom7893):

ds/dt = 7 When A = 64, s = 8 A = s^2 dA/dt = 2s*ds/dt = 2*7*8 = 2*56 = 112

OpenStudy (bahrom7893):

related rates is a nice and simple topic. U just gotta get used to it.